OpenStudy (anonymous):

What is the fourth term in this multiplication pattern? 2 x 6 n-1

OpenStudy (anonymous):

\[2^{n-1}\]

OpenStudy (anonymous):

put n=4

OpenStudy (anonymous):

@Shay17

OpenStudy (anonymous):

yes substitute 4 for n

OpenStudy (anonymous):

is that you get 432

OpenStudy (anonymous):

47

OpenStudy (anonymous):

A. 12 B. 72 C. 432 D. 2592

OpenStudy (anonymous):

6 times 2 = 12 12 times 4 = 48 48 minus 1 = 47

OpenStudy (anonymous):

Go with C. then

OpenStudy (anonymous):

idk I got what u got shay this question is weird

OpenStudy (anonymous):

idk how @matricked got 432 because he didn't explain , the #1 thing you're supposed to do

OpenStudy (anonymous):

yeah @matricked

OpenStudy (anonymous):

2*6^(4-1) =2*6^3=2*216=432

OpenStudy (anonymous):

ohh well that makes sense if it was an exponent then

OpenStudy (anonymous):

ill just go with C and hope for the best

OpenStudy (anonymous):

@Lindsay_552 please use the carrot symbol "^" if there is an exponent in your problem

OpenStudy (anonymous):

C is correct @Lindsay_552